# AndesVL Technical Report: An Efficient Mobile-side Multimodal Large Language Model

### v2

@@ Abstract
In recent years, while cloud-based MLLMs such as QwenVL, InternVL, GPT-4o, Gemini, and Claude Sonnet have demonstrated outstanding performance with enormous model sizes reaching hundreds of billions of parameters, they significantly surpass the limitations in memory, power consumption, and computing capacity of edge devices such as mobile phones. This paper introduces AndesVL, a suite of mobile-side MLLMs with 0.6B to 4B parameters based on Qwen3’s LLM and various visual encoders. We comprehensively outline the model architectures, training pipeline, and training data of AndesVL, which achieves first-tier performance across a wide range of open-source benchmarks, including fields such as text-rich image understanding, reasoning and math, multi-image comprehension, general VQA, hallucination mitigation, multilingual understanding, and GUI-related tasks when compared with state-of-the-art models of a similar scale. Furthermore, we introduce a 1+N LoRA architecture alongside a Quantization-Aware LoRA Fine-Tuning (QALFT) framework to facilitate efficient task adaptation and model compression during mobile-side deployment of AndesVL. Moreover, utilizing our cache eviction algorithm—OKV—along with customized speculative decoding and compression strategies, we achieve a 6.7x peak decoding speedup ratio, up to 30.9% memory reduction, and 1.8 bits-per-weight when deploying AndesVL-4B on MediaTek Dimensity 9500 chips. We release all models on https://huggingface.co/OPPOer.
@@ 3.3.1 Pre-train Data
Caption
Emu2 [206], ShareGPT-4V [25], Laion-ZH [190],
Wukong [64], Taisu [135]
OCR
DocMatrix [102], DocStruct [76, 77], Leopard-Instruct [83],
Pixmo-doc [46], Anyword-3M [218], PIN-14M [225],
In-house collected and synthesized (by Synthdog [225]) OCR data
Visual Grounding
Visual Genome [100], RefCOCOs [250, 125], LVIS [182],
Flickr30k-Entities [182], Groma [149]
GUI and Agent
AITW [188], AITZ [260], AMEX [19], Android Control [117], Widget Caption [118],
Rico [47], SeeClick [32], UIbert [12], Screen2Words [223], MultiUI [130],
Aria-UI[244], OS-Atlas [234], Mind2Web [48], GUI-Odyssey [144], OmniAct [92],
In-house AndesUI training set
VQA
Infinity-MM [65], MAmmoTH [70], LLaVA-OneVision [108],
The  Cauldron [105], VisualWebInstruct [84], PangeaInstruct [255]
Long CoT
OpenMathReasoning [160], OpenCodeReasoning [5], OpenThoughts [67], Nemotron [164],
In-house multiModal long CoT data
Interleaved Image-Text
MMC4 [275], MINT [11], Multimodal Textbook [264],
Wanjuan [72], OmniCorpus [116],
In-house crawled data from Chinese websites and Apps
Pure Text
Fineweb-Edu-Chinese [253], Fineweb-Edu [138], FineMath [6], OpenCoder [79],
Infinity-Instruct [175]
@@ 3.3.2 Post-train Data
VQA
EATEN [69], PMC-VQA [265], OmniAlign-V [268], Dvqa [90],
$m$-localized-narratives [183], OCR-VQA[157], Plotqa[155], ShareGPT-4o [179],
geoqa-plus[22], Figureqa [91], Tallyqa [4], LACR-I2I [165], Robut-wikisql [172],
Mimic-cgd [104], clevr [87], textvqa [199], scienceqa [141], mpdocvqa [214],
nlvr2 [203], ShareGPT4V [23], cocoqa [53], ScreenQa [75], raven [257],
unigeo [21], docmatix [103], robut-wtq [270], mapqa [20], iconqa [143], chart2text [174],
docreason25k [76], kvqa [99], scitsr [34], $m$-LADD [166], tabmwp [142],
KonIQ-10k [73], SVIT [267], sujet-finance-qa-vision [204],
viet-sharegpt-4o-text-vqa [50], ChartQA [136], cambrian-10m [216], $m$-aokvqa [191],
ai2diagram [96], Clevr-CoGenT-TrainA-70K-Complex [87], laion-gpt4v [101],
geometry3k [140], DocVQA [153], vistext [208], simchart9k [235],
geomverse [95], spot-the-diff [207], robut-sqa [270], multihiertt [269],
lrv-instruction [127], objects365 [193], mtvqa-train [209], iam [33], finqa [31],
viquae [107], fsc147-controlnet [213], $m$-tat-qa [273], infographic-vqa [63],
vsr [262], orand-car-a [60], $m$-tqa [170], $m$-intergps [169],
JourneyBench-Hallucination [88], $m$-vqarad [171], $m$-diagram-image-to-text [168],
ctw[133], naf[41], LIME-DATA-ai2d-train [124], mmc-inst [56], COCO-Text [74],
HME100K [221], st-vqa [220], fintabnet [9] , CoSyn-400K [46], PuzzleVQA [35]
OCR
anytext [218], CORD [180], invoice-to-json [1], arxiv-ocr [173], textocr [158],
invoices-and-receipts-ocr-v2 [163], mall-receipt-extraction [18],
invoices-and-receipts-ocr-v1 [162], ds-receipts-v2-train [161], dataset-receipt [80],
invoices-donut-data-v1 [94], Vision-OCR-Financial-Reports [7],
handwritten-text-ocr [215], nutritional-data-poie [93]
Pure Text Dialogue
sharegpt-gpt4 [198], ruozhiba [137], Ner-sentiment-analysis-sharegpt [246],
chinese-ner-sft [184], few-shot-ner-sft [185], SystemChat [38]
Image Captioning
Detailed Caption [122], VizWiz [71]
Code Generation
WebSight [106], $m$-datikz [167]
Function Calling
Function-Calling-Dataset-V1 [81], glaive-function-calling [62]
@@
Math
cmm-math[131], MMR1-Math-RL-Data-v0 [159], Codegebragpt-multimodal [201],
Geo170K [147], MathV360K [196], Multimath-300k [181], Unimer-math-ocr [222],
Openr1-math-220k [176], MathInstruct [254], MetaMathQA [251]
@@ 4.2.2 QALFT
To overcome this limitation, we co-developed the Quantization-Aware LoRA Fine-Tuning (QALFT) framework in collaboration with MediaTek. QALFT begins by applying PTQ to a QAT-pretrained base model and permanently freezing its quantization encodings. Subsequent LoRA weights are then trained on top of this fixed, quantized backbone—following a paradigm analogous to QLoRA [49]. This design enables independent updates of scenario-specific LoRA modules without re-quantizing the base model, thereby eliminating quantization-induced performance degradation during deployment and significantly streamlining the iteration cycle for task-specific algorithms.
@@ 5.1.2 Evaluation Results
Model Text-rich
Reasoning
&Math
Multi-image
General
VQA
Hallucination Multilingual Overall
@@ 5.2.2 Evaluation Results
Model
MMMU
MMMU
Pro
MathVista
MathVision
MathVerse
(vision-only)
DynaMath
(worst case)
WeMath LogicVista Overall
@@ 5.3.2 Evaluation Results
Model
AI2D
(w M)
AI2D
(w/o M)
ChartQA
TextVQA
DocVQA
InfoVQA
OCR
Bench
SEED
2-Plus
Overall
@@
Moreover, the advantages over existing models on text-rich tasks persist down to our smaller-scale versions. To be specific, our proposed AndesVL-2B-Instruct, AndesVL-1B-Thinking, and AndesVL-0.6B-Instruct models all rank first in their respective model-size groups, with overall scores of 82.4, 77.4, and 73.5, respectively.
@@ 5.5.2 Evaluation Results
Model MME_sum
MMBench
v1.1
MMVet MMStar RealWorldQA
R-Bench
Overall
@@
As illustrated in Table 10, the AndesVL series (4B, 1B, and 0.6B) achieve the top performance in their respective groups, while the 2B model also delivers a highly competitive result. A detailed breakdown reveals that while there remains room for improvement on the MMVet benchmark, the AndesVL series demonstrates exceptionally strong and robust performance on both MME and RealWorldQA. This suggests that our model extracts robust representations and displays a strong ability to comprehend real-world scenarios, enabling it to effectively tackle complex and dynamic tasks.
@@ 5.7.2 Evaluation Results
As demonstrated in Table 12, both the Thinking and Instruct variants of AndesVL-4B demonstrate exceptional multilingual capabilities, achieving a leading score of 64.9, which surpasses the previous best model, Ovis2-4B [145], by 1.5 points. This advantage persists in the smaller-scale variants of AndesVL, with each one achieving top multilingual scores within their respective sub-groups.
@@ 5.9 Ablation Studies
Model MMVet
MathVerse
(vision-only)
RealWorldQA OCRBench HallusionBench Overall
@@
Model
MathVista
MathVision
WeMath
MathVerse
(vision-only)
MMMU MMMU_Pro Overall
@@ 6.1 Results of Quantization-Aware Training
Model
DocVQA
InfoVQA
TextVQA
ChartQA
Overall
@@ 8 Conclusion
This paper presents AndesVL, a suite of mobile-side MLLMs with parameter sizes ranging from 0.6B to 4B. By integrating Qwen3’s LLM and various visual encoders, AndesVL achieves first-tier performance on multiple open-source benchmarks and the self-developed AndesUI benchmark, especially excelling in mobile UI understanding. The proposed 1+N LoRA architecture and Quantization-Aware LoRA Fine Tuning (QALFT) framework enable efficient task adaptation and model compression. By employing our proposed OKV, meticulously designed speculative decoding techniques and compression strategies, we can achieve 1.8 bits-per-weight, 6.7x peak decoding speed ratio and up to 30.9% memory reduction when deploying AndesVL-4B on MediaTek Dimenisity 9500 chips. This work bridges the gap between cloud-based MLLMs and edge devices, providing a practical solution for mobile-side MLLM and paving the way for future advancements in edge AI.
@@ Appendix C Qualitative Examples
In this section, a series of qualitative examples are illustrated for various capabilities of the AndesVL model by presenting responses generated from AndesVL-4B. Fig. 22 shows strong image understanding; Fig. 23, 24 and 25 demonstrate multi-image perception ability; Fig. 26 shows basic reasoning capability of AndesVL; and Fig. 27 and 28 illustrate document understanding and OCR capabilities.
Examples demonstrating AndesVL-4B-Thinking’s visual reasoning capability are provided in Fig. 29, 30 and 31. Fig. 29 shows a case of AndesVL-4B-Thinking solving a geometry problem. Fig. 31 illustrates the model’s capability of handling reasoning tasks related to charts.
UI understanding and reasoning cases of AndesVL-4B are shown in Fig. 32 and 33.

### v3

@@ Abstract
In recent years, while cloud-based MLLMs such as QwenVL, InternVL, GPT-4o, Gemini, and Claude Sonnet have demonstrated outstanding performance with enormous model sizes reaching hundreds of billions of parameters, they significantly surpass the limitations in memory, power consumption, and computing capacity of edge devices such as mobile phones. This paper introduces AndesVL, a suite of mobile-side MLLMs with 0.6B to 4B parameters based on Qwen3’s LLM and various visual encoders. We comprehensively outline the model architectures, training pipeline, and training data of AndesVL, which achieves first-tier performance across a wide range of open-source benchmarks, including fields such as text-rich image understanding, reasoning and math, multi-image comprehension, general VQA, hallucination mitigation, multilingual understanding, and GUI-related tasks when compared with state-of-the-art models of a similar scale. Furthermore, we introduce a 1+N LoRA architecture alongside a Quantization-Aware LoRA Fine-Tuning (QALFT) framework to facilitate efficient task adaptation and model compression during mobile-side deployment of AndesVL. Our QALFT experiments reveal that AndesVL maintains performance with only ignorable degradation (3%) after deployment on mobile devices compared to the original floating-point model. Moreover, utilizing our cache eviction algorithm—OKV—along with customized speculative decoding and compression strategies, we achieve a 6.7x peak decoding speedup ratio, up to 30.9% memory reduction, and 1.8 bits-per-weight when deploying AndesVL-4B on MediaTek Dimensity 9500 chips. We release all models on https://huggingface.co/OPPOer.
@@ 3.3.1 Pre-train Data
Caption Emu2 [206], ShareGPT-4V [25], Laion-ZH [190], Wukong [64], Taisu [135]
OCR DocMatrix [102], DocStruct [76, 77], Leopard-Instruct [83], Pixmo-doc [46], Anyword-3M [218], PIN-14M [225], In-house collected and synthesized (by Synthdog [225]) OCR data
Visual Grounding Visual Genome [100], RefCOCOs [250, 125], LVIS [182], Flickr30k-Entities [182], Groma [149]
GUI and Agent AITW [188], AITZ [260], AMEX [19], Android Control [117], Widget Caption [118], Rico [47], SeeClick [32], UIbert [12], Screen2Words [223], MultiUI [130], Aria-UI[244], OS-Atlas [234], Mind2Web [48], GUI-Odyssey [144], OmniAct [92], In-house AndesUI training set
VQA Infinity-MM [65], MAmmoTH [70], LLaVA-OneVision [108], The  Cauldron [105], VisualWebInstruct [84], PangeaInstruct [255]
Long CoT OpenMathReasoning [160], OpenCodeReasoning [5], OpenThoughts [67], Nemotron [164], In-house multiModal long CoT data
Interleaved Image-Text MMC4 [275], MINT [11], Multimodal Textbook [264], Wanjuan [72], OmniCorpus [116], In-house crawled data from Chinese websites and Apps
Pure Text Fineweb-Edu-Chinese [253], Fineweb-Edu [138], FineMath [6], OpenCoder [79], Infinity-Instruct [175]
@@ 3.3.2 Post-train Data
VQA EATEN [69], PMC-VQA [265], OmniAlign-V [268], Dvqa [90], mm-localized-narratives [183], OCR-VQA[157], Plotqa[155], ShareGPT-4o [179], geoqa-plus[22], Figureqa [91], Tallyqa [4], LACR-I2I [165], Robut-wikisql [172], Mimic-cgd [104], clevr [87], textvqa [199], scienceqa [141], mpdocvqa [214], nlvr2 [203], ShareGPT4V [23], cocoqa [53], ScreenQa [75], raven [257], unigeo [21], docmatix [103], robut-wtq [270], mapqa [20], iconqa [143], chart2text [174], docreason25k [76], kvqa [99], scitsr [34], mm-LADD [166], tabmwp [142], KonIQ-10k [73], SVIT [267], sujet-finance-qa-vision [204], viet-sharegpt-4o-text-vqa [50], ChartQA [136], cambrian-10m [216], mm-aokvqa [191], ai2diagram [96], Clevr-CoGenT-TrainA-70K-Complex [87], laion-gpt4v [101], geometry3k [140], DocVQA [153], vistext [208], simchart9k [235], geomverse [95], spot-the-diff [207], robut-sqa [270], multihiertt [269], lrv-instruction [127], objects365 [193], mtvqa-train [209], iam [33], finqa [31], viquae [107], fsc147-controlnet [213], mm-tat-qa [273], infographic-vqa [63], vsr [262], orand-car-a [60], mm-tqa [170], mm-intergps [169], JourneyBench-Hallucination [88], mm-vqarad [171], mm-diagram-image-to-text [168], ctw[133], naf[41], LIME-DATA-ai2d-train [124], mmc-inst [56], COCO-Text [74], HME100K [221], st-vqa [220], fintabnet [9] , CoSyn-400K [46], PuzzleVQA [35]
OCR anytext [218], CORD [180], invoice-to-json [1], arxiv-ocr [173], textocr [158], invoices-and-receipts-ocr-v2 [163], mall-receipt-extraction [18], invoices-and-receipts-ocr-v1 [162], ds-receipts-v2-train [161], dataset-receipt [80], invoices-donut-data-v1 [94], Vision-OCR-Financial-Reports [7], handwritten-text-ocr [215], nutritional-data-poie [93]
Pure Text Dialogue sharegpt-gpt4 [198], ruozhiba [137], Ner-sentiment-analysis-sharegpt [246], chinese-ner-sft [184], few-shot-ner-sft [185], SystemChat [38]
Image Captioning Detailed Caption [122], VizWiz [71]
Code Generation WebSight [106], mm-datikz [167]
Function Calling Function-Calling-Dataset-V1 [81], glaive-function-calling [62]
@@
Math cmm-math[131], MMR1-Math-RL-Data-v0 [159], Codegebragpt-multimodal [201], Geo170K [147], MathV360K [196], Multimath-300k [181], Unimer-math-ocr [222], Openr1-math-220k [176], MathInstruct [254], MetaMathQA [251]
@@ 4.2.2 QALFT
To overcome this limitation, we co-developed the Quantization-Aware LoRA Fine-Tuning (QALFT) framework in collaboration with MediaTek. QALFT begins by applying PTQ to a QAT-pretrained base model and permanently freezing its quantization encodings. Subsequent LoRA weights are then trained on top of this fixed, quantized backbone—following a paradigm analogous to QLoRA [49]. This design enables independent updates of scenario-specific LoRA modules without re-quantizing the base model, thereby eliminating quantization-induced performance degradation during deployment and significantly streamlining the iteration cycle for task-specific algorithms. Empirical evaluations demonstrate that, in this framework, the on-device performance of AndesVL degrades by only 3% relative to its full-precision model, the marginal loss validates the efficacy of QALFT in real-world applications.
@@ 5.1.2 Evaluation Results
Model Text-rich Reasoning &Math Multi-image General VQA Hallucination Multilingual Overall
@@ 5.2.2 Evaluation Results
Model MMMU (val) MMMU Pro MathVista (mini) MathVision MathVerse (vision-only) DynaMath (worst case) WeMath LogicVista Overall
@@ 5.3.2 Evaluation Results
Model AI2D (w M) AI2D (w/o M) ChartQA (test) TextVQA (val) DocVQA (test) InfoVQA (test) OCR Bench SEED 2-Plus Overall
@@
Moreover, the advantages over existing models on text-rich tasks persist down to our smaller-scale versions. To be specific, our proposed AndesVL-2B-Instruct, AndesVL-1B-Thinking, and AndesVL-0.6B-Instruct models all achieve leading results in their respective model-size groups, with overall scores of 82.4, 77.4, and 73.5, respectively.
@@ 5.5.2 Evaluation Results
Model MME_sum MMBench v1.1 MMVet MMStar RealWorldQA R-Bench (dis) Overall
@@
As illustrated in Table 10, the AndesVL series (4B, 1B, and 0.6B) achieve the top performance in their respective groups, while the 2B model also delivers a highly competitive result. This suggests that our model extracts robust representations and displays a strong ability to comprehend real-world scenarios, enabling it to effectively tackle complex and dynamic tasks.
@@ 5.7.2 Evaluation Results
As demonstrated in Table 12, both the Thinking and Instruct variants of AndesVL-4B demonstrate exceptional multilingual capabilities, achieving a leading score of 64.9, which surpasses the previous best model, Ovis2-4B [145], by 1.5 points. This advantage persists in the smaller-scale variants of AndesVL, with each one achieving leading multilingual scores within their respective sub-groups.
@@ 5.9 Ablation Studies
Model MMVet MathVerse (vision-only) RealWorldQA OCRBench HallusionBench Overall
@@
Model MathVista (mini) MathVision WeMath MathVerse (vision-only) MMMU MMMU_Pro Overall
@@ 6.1 Results of Quantization-Aware Training
Model DocVQA (test) InfoVQA (test) TextVQA (val) ChartQA (test) Overall
@@
6.2 Results of QALFT
To further improve performance across various on-device scenarios, we utilize the QLAFT framework to train LoRA weights specific to each situation. The experiments mentioned in the Table 18 are all completed based on the pre-trained AndesVL-4B-Instruct-Base model. AndesVL-4B-Instruct-Base-LoRA (floating point) represents LoRA fine-tuned floating-point model trained on OCR data, AndesVL-4B-Instruct-Base-LoRA (PTQ) represents AndesVL-4B-Instruct-Base-LoRA (floating point) with PTQ, AndesVL-4B-Instruct-Base-LoRA (QAT+PTQ) represents AndesVL-4B-Instruct-Base-LoRA (floating point) with QAT and PTQ, AndesVL-4B-Instruct-Base (QAT+PTQ+QALFT) represents QALFT training on AndesVL-4B-Instruct-Base with QAT and PTQ. The experimental results in Table 18 show that the performance of the model only with PTQ decreases significantly, QAT and QALFT can significantly improve the performance of the model on the device side, while QALFT can decrease slightly by 3% compared to the floating-point model.
LoRA Models TextVQA (val) ChartQA (test) AI2D (w M) Overall
AndesVL-4B-Instruct-Base-LoRA (floating point) 81.1 87.5 83.4 84.0
AndesVL-4B-Instruct-Base-LoRA (PTQ) 67.2 66.1 65.5 66.3
AndesVL-4B-Instruct-Base-LoRA (QAT+PTQ) 77.2 84.0 80.8 80.7
AndesVL-4B-Instruct-Base (QAT+PTQ+QALFT) 80.8 86.4 81.3 82.8
Table 18: Comparison on 3 OCR benchmarks performance among AndesVL-4B-Instruct-Base-LoRA (floating point), AndesVL-4B-Instruct-Base-LoRA (PTQ), AndesVL-4B-Instruct-Base-LoRA (QAT+PTQ) and AndesVL-4B-Instruct-Base (QAT+PTQ+QALFT), the results prove that QAT and QALFT significantly improve performance.
@@ 8 Conclusion
This paper presents AndesVL, a suite of mobile-side MLLMs with parameter sizes ranging from 0.6B to 4B. By integrating Qwen3’s LLM and various visual encoders, AndesVL achieves first-tier performance on multiple open-source benchmarks and the self-developed AndesUI benchmark, especially excelling in mobile UI understanding. The proposed 1+N LoRA architecture and Quantization-Aware LoRA Fine Tuning (QALFT) framework enable efficient task adaptation and model compression. QALFT ensures that AndesVL maintains high precision performance with only ignorable degradation (3%) after deployment on mobile devices compared to the original floating-point model. By employing our proposed OKV, meticulously designed speculative decoding techniques and compression strategies, we can achieve 1.8 bits-per-weight, 6.7x peak decoding speed ratio and up to 30.9% memory reduction when deploying AndesVL-4B on MediaTek Dimenisity 9500 chips. This work bridges the gap between cloud-based MLLMs and edge devices, providing a practical solution for mobile-side MLLM and paving the way for future advancements in edge AI.
@@ Appendix C Qualitative Examples
In this section, a series of qualitative examples are illustrated for various capabilities of the AndesVL model by presenting responses generated from AndesVL-4B. Fig. 23 shows strong image understanding; Fig. 24, 25 and 26 demonstrate multi-image perception ability; Fig. 27 shows basic reasoning capability of AndesVL; and Fig. 28 and 29 illustrate document understanding and OCR capabilities.
Examples demonstrating AndesVL-4B-Thinking’s visual reasoning capability are provided in Fig. 30, 31 and 32. Fig. 30 shows a case of AndesVL-4B-Thinking solving a geometry problem. Fig. 32 illustrates the model’s capability of handling reasoning tasks related to charts.
UI understanding and reasoning cases of AndesVL-4B are shown in Fig. 33 and 34.